OpenStudy (anonymous):

how can you tell if someone is doing work on a box? They are doing work on the box if: weighs more when they push it doesnt move backwards when they pull on it moves forward when pushed doesnt move forward when its pushed

OpenStudy (anonymous):

\[W = F \times d \times \cos \theta\] Work = force* displacement* cos theta, which means if there is no displacement with the amount of force then no work is being done. So, if the box moves forward when pushed, then there is work being done.